This is true by (citation). More stuff.

Theorem 0.1 (Wiles). Fermat’s Last Theorem is false. Not. 1. First thing.

2. second thing. Recall Theorem 0.1.

There was strange theorem on page 1. Also remember number 2 on page 1.